Find the primeter of a right triangle with legs of 10 inches and 8 inches
1
Expert's answer
2011-10-18T11:50:41-0400
using pythagorean theorem we can find the 3rd side c=sqrt(10^2+8^2)=sqrt(164)
So perimeter p=10+8+sqrt(164)=18+12.8=30.8 inches